OpenStudy (anonymous):

The average annual salary of the employees of a company in the year 2005 was $70,000. It increased by the same factor each year and in 2006, the average annual salary was $82,000. Let f(x) represent the average annual salary, in thousand dollars, after x years since 2005. Which of the following best represents the relationship between x and f(x)?

OpenStudy (anonymous):

(A) f(x) = 70(1.17)^x (B) f(x) = 82(1.17)^x (C) f(x) = 70(2.2)^x (D) f(x) = 82(2.2)^x

OpenStudy (anonymous):

Can someone plz help me with this question? @pooja195 @cwrw238 @iGreen @Michele_Laino

OpenStudy (anonymous):

@pooja195 I actually realized this one was on history. I am going to post it on mathematics.